3. This problem uses the wblake data set in the alr4 package. This data set includes samples of small mouth bass collected in West Bearskin Lake, Minnesota, in 1991. Interest is in predicting length with age. Complete this problem without using Im) in R. (a) Do the regression of length on age, and report the estimates, their standard errors and the estimate of variance. Interpret Bo and B. (b) Obtain a 90% confidence interval for A from the data. Interpret this interval. (c) Obtain a prediction and a 90% prediction interval for a small mouth bass at age 2, Interpret this interval.
Answer #1

--------------------------------------------------------------------------------------------------------------------------------------------------------------

Output:

BRStudio File Edit Code View Plots Session Build Debug Profile Tools Help * Addins Source Console Terminal # install.packages> # Part (b) solution: > # Calculate a 90% confidence interval >a=1-0.90 > t . Star = round (at ( 1-a/2, n-2 ) , 3) for beta1

Code in text format:

install.packages("alr4")
library(alr4)
data(wblake)
names(wblake)

# Part (a) solution:
X = wblake$Age; Y = wblake$Length
xbar = mean(X); ybar = mean(Y); n = length(Y)
SSxy = sum((X-xbar)*(Y-ybar))
SSxx = sum((X-xbar)^2); SSyy = sum((Y-ybar)^2)

(b1 = SSxy/SSxx); (b0 = ybar-b1*xbar) # Reg Estimates

Y.hat = b0+b1*X # Reg of length on age

SSE = sum((Y-Y.hat)^2); SSR = sum((Y.hat-ybar)^2); SST = SSR+SSE
(R.Sq = SSR/SST) # Coefficient of Deter

(MSE = SSE/(n-2)) # Estimate of Variance
(SE.b1 = sqrt(MSE/SSxx)) # SE of slope b1
(SE.b0 = sqrt(MSE*((1/n)+(xbar^2/SSxx)))) # SE of intercept b0

#------------------------------------------------------------------------
#
# Part (b) solution:
# Calculate a 90% confidence interval for beta1.
a = 1-0.90 # Confidence level = 0.90
t.star = round(qt(1-a/2,n-2),3) # t critical value
ME = t.star*SE.b1
b1-c(ME,-ME) # 90% CI for beta1

#-------------------------------------------------------------------------
#
# Part(c) solution:
# Cal a prection and a 90% preidction interval for a small mouth at age 2
x0 = 2
(Y.Pred = b0+b1*x0) # Predicted Y when X=2
ME = t.star*sqrt(MSE)*sqrt(1+(1/n)+((x0-xbar)^2)/SSxx)
Y.Pred - c(ME,-ME)